# Mixtera: A Data Plane for Foundation Model Training

### v2

@@ Abstract.
State-of-the-art large language and vision models are trained over trillions of tokens that are aggregated from a large variety of sources. As training data collections grow, manually managing the samples becomes time-consuming, tedious, and prone to errors. Yet recent research shows that the data mixture and the order in which samples are visited during training can significantly influence model accuracy. We build and present Mixtera, a data plane for foundation model training that enables users to declaratively express which data samples should be used in which proportion and in which order during training. Mixtera is a centralized, read-only layer that is deployed on top of existing training data collections and can be declaratively queried. It operates independently of the filesystem structure and supports mixtures across arbitrary properties (e.g., language, source dataset) as well as dynamic adjustment of the mixture based on model feedback. We experimentally evaluate Mixtera and show that our implementation does not bottleneck training and scales to 256 GH200 superchips. We demonstrate how Mixtera supports recent advancements in mixing strategies by implementing the proposed Adaptive Data Optimization (ADO) algorithm in the system and evaluating its performance impact. We also explore the role of mixtures for vision-language models.
PVLDB Reference Format:
PVLDB, 1(1): 1 - 1, 1.
doi:1
††This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 1, No. 1 ISSN 2150-8097.
doi:1
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/eth-easl/mixtera.
@@ 1. Introduction
Large language and vision models (often called foundation models) have become omnipresent in our daily life. They show enormous capabilities in a diverse set of tasks (Brown et al., 2020; Bommasani et al., 2022; OpenAI, 2024; Meta, 2024a; Radford et al., 2019), such as assistance with writing and coding, video understanding, and even agentic interaction with the world. The training of such language and vision models (LLMs/VLMs) presents new challenges for managing training data due to the ever-growing sizes of models and datasets. To achieve high accuracy, state-of-the-art models train over trillions of tokens from aggregated data collections such as RedPajama (Weber et al., 2024), Dolma (Soldaini et al., 2024), or FineWeb (Penedo et al., 2024a). For example, Meta’s Llama 3.3 70B model is trained on a corpus of 15 trillion tokens (Meta, 2024a, b). These collections are built based on data from various sources, such as Wikipedia or Common Crawl dumps.
Training data is typically stored on distributed filesystems in GPU clusters or data lakes in the cloud, and managed manually and ad hoc by ML engineers (Figure 2a). Selecting the right proportions out of this data with particular characteristics (e.g., language, topic, source) for training is critical for model performance (Chen et al., 2023; Ye et al., 2024; Xie et al., 2024). Individual users write ad hoc scripts to process the training data, filter relevant subsets with the properties of interest, often pre-tokenize it, and then mix it for their use case, e.g., they might want to train on 50 % data from Wikipedia and 50 % from movie subtitles. This reflects our experience working with ML researchers as part of the Swiss AI initiative which aims to develop open-source LLMs (ETH AI Center & EPFL AI Center, 2025). We confirmed through discussions with industry teams that such setups are common. This can get complex as training data may need to be mixed based on a variety of characteristics. For example, in addition to satisfying source data proportions (Wikipedia vs. movie subtitles in the previous example), we may also want the training data to be 80 % in English and 20 % in German. Recent works show that the data mixture should also be adjusted during training. The SmolLM2 model is trained with four stages of data mixtures (Allal et al., 2025). Algorithms such as Adaptive Data Optimization (ADO) (Jiang et al., 2024), Aioli (Chen et al., 2024a), PiKE (Li et al., 2025), and Skill-It (Chen et al., 2023) even propose to adjust the mixture dynamically based on the model behavior (e.g., loss per domain) during training. Figure 1 shows that using ADO can increase the performance of LLMs over a static mixture across different scales on the downstream HellaSwag benchmark (Zellers et al., 2019).
@@
Today, there is no open-source system that automatically manages vast amounts of training data based on fine-grained characteristics. Implementing data filtering and subsequent mixing requires users to manually keep track of metadata. At least in the open-source world, this is typically done as part of the directory structure of the filesystem, e.g., there is one subdirectory per source dataset, and then we sample from each directory (Figure 2a). This approach is limited because each data sample has multiple properties that can be used to determine whether it should be used for training. Filesystems fundamentally do not offer the right interface for managing training data and mixing, as they do not provide declarative query interfaces or a native way to track which model was trained on what data. Running a data processing job to fully materialize the mixed training set for each training run has lots of overhead and leads to data duplication. A streaming-oriented approach that selects, tokenizes, and mixes data on the fly provides much more flexibility and eases experimentation. Utilizing a full-fledged DBMS for tracking data properties would burden ML engineers with complexities such as database administration, schema design, and performance tuning. We need a lightweight data plane that enables users to declaratively query and mix data based on arbitrary properties, independent of the filesystem structure, and to adjust this mixture dynamically.
@@
We present Mixtera, a data plane that can be deployed on top of existing LLM/VLM training data collections. It is a centralized, read-only layer and can be declaratively queried from training clients. As shown in Figure 2b, it reduces the amount of materialization and data scripting necessary for training jobs. While existing data loaders as outlined in Table 1 are limited by filesystem constraints, Mixtera supports arbitrary properties independent of the filesystem, and makes it easy for model engineers to experiment with different filter criteria, fixed learning curricula, and fully dynamic mixing algorithms that learn the mixture during training. Mixtera follows a client-server model. For a training job, the server, which indexes and manages the samples, first statically filters out the relevant samples, and then continuously distributes chunks to the clients. Chunks are fixed-size collections of pointers to samples in files, and adhere to the current mixture. The clients then fetch the relevant samples and return the relevant data to the training loop. We design and implement Mixtera tailored to the needs for foundation model training, and make the following key contributions:
Mixtera enables users to declaratively specify mixtures across properties independent of the filesystem structure and training frameworks by indexing all samples and their properties.
Mixtera enables dynamically changing the data mixture and the properties determining the mixture during training. It achieves this by generating and streaming chunks, i.e., fixed-size lists of pointers to samples following the current mixture.
We show Mixtera does not bottleneck training and is versatile, enabling model accuracy improvements for text-only and text-vision models. As an example, we demonstrate how to implement the ADO dynamic data mixing algorithm in Mixtera and its positive impact on model accuracy.
@@ 2. Background
Foundation models are large-scale deep learning models suitable for a variety of tasks (Bommasani et al., 2022; Competition and Markets Authority, 2013). We focus on text-generation models, i.e., autoregressive large language models (LLMs) and multimodal vision-language models (VLMs). As of 2025, most such models are based on the Transformer architecture (Vaswani et al., 2017). They are trained on huge corpora of training data in a self-supervised manner to maximize the likelihood of predicting the tokens of a training sequence.
Training phases. Training is structured into pre-training and post-training phases. In pre-training, we train a randomly initialized model on a general purpose data corpus (Section 2.1) to derive a base model. In post-training, common steps include supervised finetuning (SFT) and alignment. The data used in SFT is human-selected for a specific task. In this paper, we focus on the pre-training use case.
Distributed training. Training foundation models requires distributing computation across multiple GPUs. Training frameworks typically employ 3D parallelism (Ben-Nun and Hoefler, 2019; Hoefler et al., 2024), consisting of pipeline parallelism (PP), i.e., partioning the model layers between devices (Narayanan et al., 2021; Huang et al., 2019; Narayanan et al., 2019), tensor parallelism (TP), i.e., splitting individual tensor operations within layers across devices (Shazeer et al., 2018; Shoeybi et al., 2019), and data parallelism (DP), i.e., replication of the model across device groups. PP and TP together are referred to as model parallelism. Nodes within the same DP group process identical inputs, while nodes across DP groups receive different data. As an extension to DP, fully-sharded data parallelism (FSDP) shards model parameters, gradients, and optimizer states across data-parallel workers (Zhao et al., 2023).
@@ 2.1. Training Data and Data Mixing
The data used for pre-training stems from data collections that include data from various sources, such as Wikipedia, Common Crawl dumps, or arXiv papers. Public examples of such collections include RedPajama (Weber et al., 2024), Dolma (Soldaini et al., 2024), and FineWeb (Penedo et al., 2024a). Besides the aggregation of data from different sources, data engineers typically clean the data, which typically involves deduplicating, filtering (e.g., the removal of personal identifiable information - PII), and applying classifiers to the data samples (e.g., to obtain a toxicity score for each sample). Longpre et al. (Longpre et al., 2024) and Penedo et al. (Penedo et al., 2024a) provide a great overview of the processing steps and their impact.
Data properties and mixtures. Each sample in the resulting collection has properties, such as its source (e.g., Wikipedia) or its language (e.g., English). Beyond the filtering operations, ML engineers need to define a data mixture. A mixture describes how the data is mixed based on their characteristics, i.e., we might train on 50 % data from Common Crawl and 50 % from movie subtitles. The data can be combined based on multiple characteristics simultaneously. For instance, besides Common Crawl and movie subtitles, we might also use 80 % French and 20 % Italian data. The granularity on which the mixture is ensured depends on each training setup and there is no common agreement. For example, a mixture could be ensured within a batch, or across a window of several batches.
Mixing algorithms. Selecting the best mixture is critical for model performance (Chen et al., 2023; Ye et al., 2024; Xie et al., 2024; Shen et al., 2024). We differentiate static mixtures, i.e., mixtures that remain constant over the entire training job, and dynamic mixtures, i.e., mixtures that change during the training job. In the last years, several algorithms for finding the best static mixture or how to adjust the mixture dynamically during training have been proposed. Algorithms such as DoReMi (Xie et al., 2024) or the data mixing laws (Ye et al., 2024) find the a static mixture via small proxy models.
Curriculum learning is an example of a pre-defined dynamic mixture. Xu et al. (Xu et al., 2024b) order samples from easy to hard to improve alignment. The SmolLM2 model was trained on 4 stages of mixtures (Allal et al., 2025). Multilingual models are often first trained on English data, and then data from other languages is included (Richburg and Carpuat, 2024; Xu et al., 2024a).
Beyond such pre-defined schedules, there is also work on adapting the data mixture to the model training dynamics, e.g., by increasing the weight of data domains which have high loss. Albalak et al. (Albalak et al., 2023) use a multi-armed bandit strategy to adjust the mixture. Skill-it orders “skills” based on model feedback (Chen et al., 2023). Aioli builds upon Skill-it and provides a unified framework for estimating the best mixture during training (Chen et al., 2024a). PiKE relies on gradient interactions (Li et al., 2025). In this paper, we use Adaptive Data Optimization (ADO) (Jiang et al., 2024) as an example of a dynamic mixing algorithms.
@@ 2.1.1. Adaptive Data Optimization
Adaptive Data Optimization (ADO) is a dynamic mixing algorithm that adjusts the data mixture during training based on the model’s learning progress on each domain (Jiang et al., 2024). The key idea is to prioritize domains where the model shows rapid improvement while considering how much each domain contributes to its own progress. ADO uses neural scaling laws to model how the loss of each domain decreases with the number of training samples . To this end, it fits a power law for each domain. Here, represents the irreducible loss of the domain, is a scaling factor, and determines how quickly the loss decreases. The parameters are re-fitted during training. The algorithm combines two components to determine the mixture weights. First, it estimates the learning speed for each domain using the derivative of the scaling law. Second, ADO maintains a credit assignment score that indicates how much each domain contributes to its own progress, based on its recent sampling frequency. These components are combined with a prior (initial) distribution to compute an intermediate preference distribution . To ensure stability, the final distribution is then computed as a weighted average between and ’s temporal average. Additionally, ADO enforces a minimum sampling probability for each domain. For more details, we refer to Jiang et al. (Jiang et al., 2024).
@@ 3. Current Challenges
Challenge 1: High engineering effort for data preparation. The current approach to preparing training data sets involves many manual offline steps with general-purpose data processing and scripting frameworks (Figure 2a). For the offline cleaning step (Section 2.1), ML engineers typically leverage data processing frameworks like Spark (Zaharia et al., 2016), Beam (Akidau et al., 2015), Data-Juicer (Chen et al., 2024b), or datatrove (Penedo et al., 2024b).
The subsequent data mixing can happen offline or online. If data is mixed offline, engineers write ad hoc scripts which create a new mixed copy of the cleaned dataset for each training run. This makes it difficult to get a quick sense of how a data mixture will impact model training when exploring different mixing policies. Some existing data loaders support online data mixing. However, they only implement it across directories, i.e., they assign a weight to each directory and sample data according to the weights. Hence, the directories need to reflect the property we want to mix on, e.g., one directory per language or data source. This inflexible approach neither supports switching the property we mix on nor supports specifying hierarchical mixtures across arbitrary properties, e.g., specify a proportion between Wikipedia and movie subtitles, and between English and German (c.f. Table 1).
Challenge 2: Inefficient data management on filesystems. Training data is typically stored and managed as files without a proper data management system, which can lead to storage overhead, performance bottlenecks, and consistency issues. Both the source data and mixed training set are typically stored in formats such as jsonl or parquet files in a shared, distributed filesystem. Existing data loaders just wrap around the filesystem and are therefore limited by its constraints (Table 1). Filesystems are commonly used because current database management systems are not natively built for foundation model training data (Wang et al., 2023). Using a proper DBMS for such data would require ML developers to define table schemas and to orchestrate dataflow from the DBMS to the model training, in addition to administration overhead to operate the DBMS itself. This situation is problematic, as (i) both the metadata and actual payloads are commonly duplicated across training jobs, (ii) filesystems do not provide an efficient interfaces to query data, (iii) there is no native way of tracking which model was trained on which data if it is accessed via general-purpose filesystem calls. Some frameworks like Megatron (Shoeybi et al., 2019; Narayanan et al., 2021) even require pre-tokenized data, which leads to duplication of the even bigger tokenized data files.
@@
Challenge 3: Rapidly evolving research. How to train the best model on a given dataset is an active area of research, with many new techniques such as dynamic mixing emerging. Offline preparation of the mixture or online mixing based on fixed directory weights does not support dynamic mixture at all. Even for researchers who are familiar with the latest mixing techniques, implementing modern mixing algorithms in a training pipeline is a painful, tedious, and error-prone task. The codebases for mixing algorithms are often tailored directly to the training framework, as well as the data collection and properties used in the respective papers. This hinders the adoption of new methods and makes researching, reproducing, and comparing different strategies difficult.

### v3

@@ Abstract.
State-of-the-art large language and vision models are trained over trillions of tokens that are aggregated from a large variety of sources. As training data collections grow, manually managing the samples becomes time-consuming, tedious, and prone to errors. Yet recent research shows that the data mixture and the order in which samples are visited during training can significantly influence model accuracy. We build and present Mixtera, a data plane for foundation model training that enables users to declaratively express which data samples should be used in which proportion and in which order during training. Mixtera is a centralized, read-only layer that is deployed on top of existing training data collections and can be declaratively queried. It operates independently of the filesystem structure and supports mixtures across arbitrary properties (e.g., language, source dataset) as well as dynamic adjustment of the mixture based on model feedback. We experimentally evaluate Mixtera and show that our implementation does not bottleneck training and scales to 256 GH200 superchips. We demonstrate how Mixtera supports recent advancements in mixing strategies by implementing the Adaptive Data Optimization (ADO) algorithm in the system and evaluating its performance impact. We also show how Mixtera enables exploring the role of mixtures for vision-language models, which is a growing area of research.
@@ 1. Introduction
Large language and vision models (LLMs/VLMs, often called foundation models) have become omnipresent in our daily lives. They show enormous capabilities in a diverse set of tasks (Brown et al., 2020; Bommasani et al., 2022; OpenAI, 2024; Meta, 2024a; Radford et al., 2019), such as assistance with writing and coding, video understanding, and even agentic interaction with the world. The training of such language and vision models presents new challenges for managing training data due to the ever-growing sizes of models and datasets. To achieve high accuracy, state-of-the-art models train over trillions of tokens. For example, Meta’s Llama 3.3 70B model is trained on a corpus of 15 trillion tokens (Meta, 2024a, b). These tokens typically come from aggregated data collections such as RedPajama (Weber et al., 2024), Dolma (Soldaini et al., 2024), or FineWeb (Penedo et al., 2024a), which include data from various sources, such as Wikipedia or Common Crawl dumps.
The composition of training data is critical to model quality (Chen et al., 2024b). Hence, selecting the right proportions of data with particular characteristics (e.g., language, topic, source) has become an active area of research to improve model performance without increasing training compute budget (Chen et al., 2023; Ye et al., 2024; Xie et al., 2024). For example, Hugging Face’s SmolLM2 model is trained with four stages of data mixtures that combine web text with specialized math, code, and instruction-following data in varying proportions (Allal et al., 2025). Algorithms such as Adaptive Data Optimization (ADO) (Jiang et al., 2024), Aioli (Chen et al., 2024a), PiKE (Li et al., 2025), and Skill-It (Chen et al., 2023) even propose adjusting the data mixture dynamically based on the model behavior (e.g., loss per domain) during training. Figure˜1 shows that ADO increases the accuracy of 1.6B and 3.6B Llama-models (c.f. Table˜2) compared to using a static mixture on the downstream HellaSwag benchmark (Zellers et al., 2019).
However, the process of composing training data mixtures today is manual, ad hoc, and error-prone (Figure˜2a). Training data is typically stored on distributed filesystems in GPU clusters or data lakes in the cloud. ML engineers and researchers write ad hoc scripts to process the training data, filter relevant subsets with the properties of interest, often pre-tokenize it, and then mix it for their use case. For example, a model developer may want to train on 50 % data from Wikipedia and 50 % from movie subtitles. This can quickly become more complex as training data may need to be mixed based on multiple characteristics. For example, in addition to source data proportions (Wikipedia vs. movie subtitles), the developer may also want the training data to be 80 % in English and 20 % in German. When using mixing algorithms such as ADO, the data mixture may also need to be adjusted on the fly during training.
@@
Model developers currently lack an open-source solution to efficiently manage and declaratively query vast amounts of training data based on the characteristics of individual data elements. This reflects our experience working with ML researchers as part of a large-scale initiative developing open-source LLMs and our conversations with industry teams. Implementing data filtering and subsequent mixing requires today’s model developers to manually keep track of metadata. At least in the open-source world, developers typically implement this as part of the directory structure of the filesystem, e.g., developers create one subdirectory per source dataset and then sample from each directory (Figure˜2a). This approach is limited because each data sample has multiple properties that can be used to determine whether it should be used for training. Filesystems fundamentally do not offer the right interface for managing training data and mixing, as they do not provide declarative query interfaces or a native way to track which model was trained on what data. Running an offline data processing job with frameworks like Apache Beam or Spark to fully materialize the mixed training set for each training run makes it difficult to quickly iterate on different mixtures during the exploration phase and leads to data duplication, increasing storage costs. While databases offer a declarative query interface, using a fully-fledged DBMS to track data properties would burden ML engineers with database administration, schema design, and database performance tuning. Having a simple read-only SQLite table at each training client might seem like a straightforward solution, but it cannot dynamically adjust data mixtures and it does not interface with training frameworks like Torchtitan (Liang et al., 2024). We need a flexible data plane that lets users easily find and combine data based on any characteristic they choose, without being restricted by how files are organized. This system should also allow users to change the data mixture on the fly and work seamlessly with current training frameworks.
@@
We present Mixtera, a data plane that can be deployed on top of existing LLM/VLM training data collections stored in distributed filesystems or cloud data lakes. Mixtera follows a client-server model. The user indexes all sample metadata at the server once. During a training job, the server statically filters out the relevant samples using SPJ-style predicates, and then continuously distributes chunks to the clients. Chunks are fixed-size collections of pointers to samples (Figure˜2b) in files that adhere to the current mixture. The clients fetch the relevant samples and return the relevant data to the training loop. By storing metadata, and deferring reads to the clients, Mixtera scales to real-world datasets. Chunks also enable dynamic data mixing and avoid re-materializing mixed data. This approach also improves the iteration speed for model developers. While existing data loaders, as outlined in Table˜1, rely on sampling from filesystem directories, Mixtera supports arbitrary properties independent of the filesystem, and makes it easy for model engineers to experiment with different filter criteria, fixed learning curricula (Allal et al., 2025; Xu et al., 2024b), and fully dynamic mixing algorithms that learn the mixture during training (Jiang et al., 2024; Chen et al., 2023, 2024a). We design and implement Mixtera tailored to the needs of foundation model training, and contribute the following:
Mixtera indexes all samples and their properties, enabling users to declaratively specify mixtures across properties independent of the filesystem structure and training frameworks.
Mixtera enables dynamically changing the data mixture and the properties used for mixing. It achieves this by generating and streaming chunks, i.e., fixed-size lists of pointers to samples following the current mixture.
Mixtera’s data fetching scales to meet the ingestion throughput demands of large-scale training jobs. We run benchmarks spanning 256 GH200 superchips and show that Mixtera does not limit training throughput. As an example for how it enables model accuracy improvements, we demonstrate how to implement the ADO dynamic data mixing algorithm in Mixtera and its positive impact on model accuracy.
@@ 2. Background
Foundation models are large-scale deep learning models suitable for a variety of tasks (Bommasani et al., 2022; Competition and Markets Authority, 2013). We focus on text-generation models, i.e., autoregressive large language models (LLMs) and multimodal vision-language models (VLMs). As of 2025, most such models are based on the Transformer architecture (Vaswani et al., 2017). They are trained on vast corpora of training data in a self-supervised manner to maximize the likelihood of predicting the tokens of a training sequence.
Training phases. Training is structured into pre-training and post-training phases. In pre-training, we train a randomly initialized model on a general-purpose data corpus (Section˜2.1) to derive a base model. In post-training, common steps include supervised finetuning (SFT) and alignment. In this paper, we perform only pre-training experiments due to space constraints, but Mixtera can also be used for post-training.
Distributed training. Training foundation models requires distributing computation across multiple GPUs. Training frameworks typically employ 3D parallelism (Ben-Nun and Hoefler, 2019; Hoefler et al., 2024), consisting of pipeline parallelism (PP), i.e., partitioning the model layers between devices (Narayanan et al., 2021; Huang et al., 2019; Narayanan et al., 2019), tensor parallelism (TP), i.e., splitting individual tensor operations within layers across devices (Shazeer et al., 2018; Shoeybi et al., 2019), and data parallelism (DP), i.e., replication of the model across device groups. PP and TP together are referred to as model parallelism. Nodes within the same DP group process identical inputs, while nodes across DP groups receive different data. As an extension to DP, fully-sharded data parallelism (FSDP) shards model parameters, gradients, and optimizer states across data-parallel workers (Zhao et al., 2023).
@@ 2.1. Training Data and Data Mixing
Pre-training data stems from data collections that include samples from various sources (e.g., Wikipedia, Common Crawl dumps, or arXiv papers). Public examples of such collections include RedPajama (Weber et al., 2024), Dolma (Soldaini et al., 2024), and FineWeb (Penedo et al., 2024a). Besides aggregating data from different sources, data engineers typically clean the data, which usually involves deduplicating, filtering (e.g., removing personal identifiable information), and applying classifiers to the data samples (e.g., to obtain a toxicity score for each sample) (Longpre et al., 2024; Penedo et al., 2024a).
Data properties and mixtures. Each data sample has properties, such as its source (e.g., Wikipedia) or its language (e.g., English). ML engineers need to define a data mixture, which describes how the data is combined based on its characteristics, e.g., we can train on 50 % data from Common Crawl and 50 % from movie subtitles. The data can be combined based on multiple characteristics simultaneously. For instance, besides Common Crawl and movie subtitles, we might also use 80 % French and 20 % Italian data.
Mixing algorithms. Selecting the best mixture is critical for model performance (Chen et al., 2023; Ye et al., 2024; Xie et al., 2024; Shen et al., 2024). We differentiate static mixtures, i.e., mixtures that remain constant over the entire training job, and dynamic mixtures, i.e., mixtures that change during the training job. Recent research proposes several algorithms for finding the best static mixture or how to adjust the mixture dynamically during training. Algorithms such as DoReMi (Xie et al., 2024) or the data mixing laws (Ye et al., 2024) find a static mixture via small proxy models.
Curriculum learning is an example of a pre-defined dynamic mixture. Xu et al. (Xu et al., 2024b) order samples from easy to hard to improve alignment. The SmolLM2 model was trained on 4 stages of mixtures (Allal et al., 2025). Multilingual models are often first trained on English data, followed by samples from other languages (Richburg and Carpuat, 2024; Xu et al., 2024a).
Beyond such pre-defined schedules, for text-only models, there is also work on adapting the data mixture to the model training dynamics, e.g., by increasing the weight of data domains that have high loss. Albalak et al. (Albalak et al., 2023) model mixture components as arms of a multi-armed bandit. Skill-it orders “skills” based on model feedback (Chen et al., 2023). Aioli builds upon Skill-it and provides a unified framework for estimating the best mixture during training (Chen et al., 2024a). PiKE relies on gradient interactions (Li et al., 2025). In this paper, we use Adaptive Data Optimization (ADO) (Jiang et al., 2024) as an example of a dynamic mixing algorithm.
@@ 2.1.1. Adaptive Data Optimization
Adaptive Data Optimization (ADO) is a dynamic mixing algorithm that adjusts the data mixture during training based on the model’s learning progress on each domain (Jiang et al., 2024). The key idea is to prioritize domains where the model shows rapid improvement while considering how much each domain benefits from its own samples. ADO uses neural scaling laws to model how the loss of each domain decreases with the number of training samples . To this end, it fits a power law for each domain. Here, represents the irreducible loss of the domain, is a scaling factor, and determines how quickly the loss decreases. The parameters are re-fitted during training. The algorithm combines two components to determine the mixture weights. First, it estimates the learning speed for each domain using the derivative of the scaling law. Second, ADO maintains a credit assignment score that indicates how much each domain contributes to its own progress, based on its recent sampling frequency. These components are combined with a prior (initial) distribution to compute an intermediate preference distribution . To ensure stability, the final distribution is then computed as a weighted average between and ’s temporal average. Additionally, ADO enforces a minimum sampling probability for each domain.
@@ 3. Current Challenges
Challenge 1: Today’s training data storage systems lack expressive, declarative APIs for data mixing. Training data is typically stored and managed as files on distributed filesystems or objects on cloud storage. As these systems are not natively built for foundation model training data (Wang et al., 2023), they lack operators for data selection, mixing, and ingestion into a training framework. This also complicates lineage tracking, as there is no native way of tracking which model was trained on which data when the data is accessed via general-purpose filesystem calls.
Challenge 2: Preparing and materializing data mixtures with offline preprocessing limits flexibility and increases storage costs. The current approach to preparing training data involves numerous manual offline steps with general-purpose data processing and scripting frameworks (Figure˜2a). For the offline cleaning step (Section˜2.1), ML engineers typically leverage data processing frameworks like Spark (Zaharia et al., 2016), Beam (Akidau et al., 2015), Data-Juicer (Chen et al., 2024b), or datatrove (Penedo et al., 2024b). Subsequent data mixing can happen offline or online. Current online data loaders sample data from directories that reflect the mixing property, e.g., one directory per source dataset (Table˜1), and therefore do not support switching the property we mix on nor specifying hierarchical mixtures across arbitrary properties. Due to the limited functionality of current online solutions, engineers often write ad hoc offline mixing scripts that create a new mixed copy of the cleaned dataset for each training run, which can drastically increase storage usage and cost. As some frameworks like Megatron (Shoeybi et al., 2019; Narayanan et al., 2021) even require pre-tokenized data, this leads to duplication of the even bigger tokenized data files. Offline mixing also drastically increases the iteration time and makes it difficult to get a quick sense of how a mixture will impact model training when exploring different mixing policies.
@@
Challenge 3: Lack of support for dynamic mixtures during training. How to train the best model on a given dataset is an active research area, with dynamic mixing emerging as a prominent new technique. Offline preparation of the mixture or other approaches such as online mixing based on fixed directory weights or using a vanilla DBMS without additional infrastructure does not support dynamic mixture at all. Additionally, even for researchers who are familiar with the latest mixing techniques, implementing modern mixing algorithms in a training pipeline is a painful, tedious, and error-prone task, as the codebases for mixing algorithms are often tailored directly to the training framework, as well as the data collection and properties used in the respective papers. This hinders the adoption of dynamic mixing algorithms and makes researching, reproducing, and comparing difficult.